OpenStudy (anonymous):

For the limit below illustrate the definition by finding the largest possible values of δ that correspond to ε = 0.2 and ε = 0.1 lim x->2 (x^3-4x+1)=1 if ε = 0.2 the largest possible value for δ is: if ε = 0.1 the largest possible value for δ is: You don't have to solve the problem, just walking me through how to do the problem step by step would help immensely. Thanks!

OpenStudy (freckles):

Did you try using the epsilon delta proof:

OpenStudy (freckles):

definition*

OpenStudy (freckles):

\[\lim_{x \rightarrow c}f(x)=L\] says: \[\text{ For all } \epsilon>0 \text{ there exists } \delta >0 \text{ such that} \\ \text{ for all } \text{ } x \text{ } 0<|x-c|<\delta => |f(x)-L|<\epsilon\] So if we chose epsilon to be 0.01 and f(x)=x^2-x and c=1. Then L=0. So we have \[ \text{ there exists } \delta >0 \text{ such taht } \\ \text{ for all } x \text{ } 0<|x-1|<\delta \\ |f(x)-L|=|x^2-x-0|<0.01\] .... Or try to look at this: Solve both of the equations for x: \[f(x)=L+\epsilon\] \[f(x)=L-\epsilon\] Then choose the x that is closest to x=1. Then plug that x into |x-c|

OpenStudy (freckles):

So we have \[x^2-x=.01\] x=1.009 x=-0.00990195 \[x^2-x=-.01\] x=.98989898 x=.0101021 The x that is closest to our c=1 is.... x=1.009 So we have \[0<|x-1|<\delta\] So plug in x=1.009 |1.009-1|=|.009|=.009 We will choose that number to be our delta... We still should be able to close by letting delta be between 0 and 1/2 So we have 0<|x-1|<delta if we choose delta to be between 0 and 1/2 we have |x-1|<1/2 -1/2<x-1<1/2 1/2<x<3/2 So we also have \[ |f(x)-L|=|x^2-x-0| =|x|*|x-1|<\epsilon\] So This means \[|x-1|<\frac{\epsilon}{|x|}\] x is between 1/2 and 3/2 So \[|x-1|<\frac{ 2 \epsilon}{3}\] \[|x-1|<\frac{2}{3}(.01)=\frac{1}{150}\] but as you see the delta from before was larger than this one so that method above might be better

OpenStudy (anonymous):

A slightly different (but fundamentally similar) approach would be to prove the limit using the definition, then plugging in the given \(\epsilon\). \[\begin{align*}\left|x^3-4x+1-1\right|&=\left|x^3-4x\right|\\\\ &=|x||x+2||x-2|\\\\ &<3\times5~|x-2|&\text{(agree to set }\delta\le1)\\\\ &<\epsilon \end{align*}\] which would mean \(\delta=\min\left\{1,\dfrac{\epsilon}{15}\right\}\). For \(\epsilon=0.2\), you have \(\delta=\dfrac{0.2}{15}=\dfrac{2}{150}\), and for \(\epsilon=0.1\), \(\delta=\dfrac{1}{150}\) (as freckles has shown).

OpenStudy (freckles):

your delta is 2/150=1/75

OpenStudy (freckles):

oops nevermind

OpenStudy (freckles):

looking at the other epsilon and didn't look anything in between

OpenStudy (freckles):

I wonder if that method gives us the largest delta He is looking for largest delta

OpenStudy (freckles):

I think you have to find when f(x)=L+epsilon and find when f(x)=L-epsilon and then get the x from there that is closest to your c (then number x is approaching) Then take that c and evaluate |x-c| to find the largest possible delta

OpenStudy (freckles):

could be wrong though